# Unraveling additional quantum many-body scars of the spin-$1$ XYXY model with Fock-space cages and commutant algebras

### v1

@@ IV.4 Additional family: Mirror-dimer states with unconstrained spins
where each term couples two nearest-neighbors that are mirror images with respect to the chosen axis. Again the Hamitonian $H_{XY}{\equiv}\sum_{j}\mathcal{M}_{j}^{k}$. We find that $H_{XY}$ possesses a set of simultaneous zero modes of all the $\mathcal{M}_{j}^{k}$, which we denote by $|\mathbb{M}_{m,m^{\prime}}^{k}\rangle$ (see Appendix C.4):
@@
where $m,m^{\prime}{\in}\{-1,0,1\}$ and $|\phi^{\pm}\rangle$ are maximally entangled spin-$1$ dimers defined previously in Eq. (31). Each state $|\mathbb{M}_{m,m^{\prime}}^{k}\rangle$ therefore comprises a fully dimerized background symmetric about the reflection axis through $(k,k+L)$, along with two freely configurable spins at the ends of this axis in the state $|m,m^{\prime}\rangle$, which can take any of the nine possible two-site spin-$1$ configurations. The total magnetization of the state $|\mathbb{M}_{m,m^{\prime}}^{k}\rangle$ is $M_{m,m^{\prime}}{=}m{+}m^{\prime}$, allowing these states to appear in both even- and odd-magnetization sectors. This does not contradict the earlier chiral-symmetry arguments, which showed that the number of zero-energy eigenstates is lower bounded by 0 in the odd-magnetization sectors [see Eq. (6)]; states can certainly exist even if this lower bound is zero. A representative state $|\mathbb{M}_{m,m^{\prime}}^{1}\rangle$ for 2L=122L{=}12 is shown in Fig. 12(b).
Furthermore, these mirror-dimer states $|\mathbb{M}_{m,m^{\prime}}^{k}\rangle$ are also annihilated by the NNNN $XY$ interaction term $V_{\rm NNNN}$ defined in Eq. (26), making them exact eigenstate of a fully non-integrable Hamiltonian $H_{\rm mirror}{=}H_{0}{+}V_{\rm NNNN}$ with energy determined by magnetization as $E_{m,m^{\prime}}^{k}{=}h(m{+}m^{\prime})$. The entanglement properties of these states depend strongly on the bipartition, as expected from their structure. If the subsystem lies entirely on one side of the reflection axis, i.e., contains exactly one site from each entangled dimer, the EE grows linearly with the subsystem size, displaying volume-law scaling as each dimer cut by the bipartition contributes a fixed entropy of $\ln(3)$. In contrast, if the subsystem includes only mirror-reflected site pairs, i.e., both sites of the dimers reside in the same subsystem, the state factorizes and its entanglement entropy vanishes.
@@ V Conclusion and outlook
Partly motivated by recent works on identifying non-thermal eigenstates from quantum interference effects in Fock space, i.e., so-called Fock-Space Cage (FSC) states [72, 73, 74], we studied the Fock-space graph of this model to uncover an entire hierarchy of FSC eigenstates, parameterized by an integer separation $r$ (Sec. III), each realizing a distinct interference-protected cage structure within the zero-energy manifold. We used this perspective to also show that previously known scars [32, 33] also admit a natural interpretation as FSC states, where destructive interference confines their support to closed loops in the Fock-space graph. These novel quasiparticle towers of eigenstates we found are also isolated from the thermal continuum in entanglement and exhibit subextensive entanglement entropy, appearing as clear outliers below the Page value. Coherent superpositions of these ladder states-while not simple product states but possessing finite, low entanglement even in the thermodynamic limit-display long-lived fidelity oscillations, the defining dynamical hallmark of scarring.
@@
The lowest members of these families are simple-looking on the Fock-space-cage graph and admit transparent geometric interpretations based on simple pairwise interference cancellations, which qualify them as FSC states. However, for higher members, the interference patterns become increasingly complex and multipath in nature, making the geometric picture less intuitive and pointing toward a deeper organizing principle. To uncover this hidden principle, we turned to ideas from the commutant algebra framework (Sec. IV), which motivates the understanding of QMBS as simultaneous eigenstates of multiple non-commuting local operators.
@@
By systematically reorganizing the terms in the spin-$1$ XYXY Hamiltonian into carefully chosen families of non-commuting operators—constructed from periodically clustered, antipodal, and mirror-symmetric combinations of exchange bonds—we identified the algebraic structure that leads to the existence of the FSC states, and also uncovered two additional classes of exact eigenstates. The first of these, arising from antipodal magnon pairings, displays a striking bipartition-dependent entanglement structure: while exhibiting volume-law scaling under standard cuts, they show sub-extensive EE under fine-tuned antipodal partitions. The second family, consisting of mirror-dimer states, features reflection-symmetric dimer configurations with two unconstrained spins. These unconstrained spins are completely free—any choice of their local configuration yields an exact zero-energy eigenstate of $H_{XY}$, which hints at their potential applications in quantum information processing.
@@
A concise summary of all exact scar families, both previously known and newly discovered, is presented in Table 2. The table lists their total magnetizations, minimal generating operator sets within the commutant algebra framework, and characteristic entanglement properties, serving as a unified map of how many kinds of nonthermal manifolds of states — from frustration-free bimagnons to highly entangled mirror-dimer states—emerge from common algebraic principles. In the future, it would be important to uncover other associated properties of these states, such as the associated asymptotic quantum many-body scars [101, 102, 103, 104, 105] and their dynamical signatures, and potentially a complete classification of the distinct kinds of local parent Hamiltonians that could realize these as eigenstates [35, 95, 77, 106]. Finally, it would also be interesting to explore whether there are yet additional QMBS hiding in the spin-$1$ XYXY model that might be revealed as simultaneous eigenstates of other sets of non-commuting operators that we have not considered in this work.
@@ Acknowledgements.
This work was undertaken on the Nandadevi supercomputer, which is maintained and supported by the High-Performance Computing Center at the Institute of Mathematical Sciences, India. S. Mohapatra is grateful to the School of Natural Sciences at the Technical University of Munich (TUM) for its hospitality in the summer of 2025. S. Moudgalya acknowledges support from the Munich Center for Quantum Science and Technology (MCQST) and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC–2111–390814868.
@@ C.3.1 Absence of a single raising operator for the volume-entangled tower
$|\psi_{\rm vol}^{\rm init}\rangle{=}\frac{1}{\sqrt{|\mathcal{D}_{L}|}}\sum_{|\vec{b^{\prime}}\rangle{\in}\mathcal{D}_{L}}\hat{\mathcal{C}}|\vec{b^{\prime}}\rangle_{1,{\cdots},L}{\otimes}|\vec{b^{\prime}}\rangle_{L{+}1,{\cdots},2L},$ (123)
@@
$|\mathbb{V}_{n}\rangle{=}\frac{1}{\sqrt{|\mathcal{D}_{L}^{M_{n}/2}|}}\sum_{|\vec{b}\rangle{\in}\mathcal{D}_{L}^{M_{n}/2}}\hat{\mathcal{C}}|\vec{b}\rangle_{1,{\cdots},L}{\otimes}|\vec{b}\rangle_{L{+}1,{\cdots},2L},$ (124)
where $\vec{b}$ runs over all spin-$1$ basis configurations with magnetization $M_{n}/2$ and $\mathcal{D}_{L}^{M_{n}/2}$ denote this Hilbert space dimension. Eq. (124) represents the Schmidt decomposition of $|\mathbb{V}_{n}\rangle$ for the standard bipartition $A_{s}{=}\{1,{\cdots},L\},~~B_{s}{=}\{L{+}1,{\cdots},2L\}$) with $|\mathcal{D}_{L}^{M_{n}/2}|$, with all non-zero Schmidt coefficients, $\lambda_{k}$, equal to $1/{\sqrt{|\mathcal{D}_{L}^{M_{n}/2}|}}$. Thus, the von Neumann EE of $|\mathbb{V}_{n}\rangle$ for this bipartition is
$\mathcal{S}_{A_{s}}(n){=}\sum_{k}\lambda_{k}^{2}\ln\lambda_{k}^{2}{=}\ln(|\mathcal{D}_{L}^{M_{n}/2}|).$ (125)
We now evaluate the $\mathcal{S}_{A_{s}}(n{=}L)$ for the mid-spectrum state $|\mathbb{V}_{n{=}L}\rangle$ (lying in the magnetization sector $M_{L}{=}0)$ in the large $L$ limit. We recall that the trinomial coefficient $|\mathcal{D}_{a}^{b}|$ (as defined in Eq. (5)) represents the coefficient of $x^{b}$ in the expansion of $(1{+}x{+}x^{{-}1})^{a}$. For large values of $a$, following a similar derivation as done in App. A, we get the approximate Gaussian form of $|\mathcal{D}_{a}^{b}|$:
$|\mathcal{D}_{a}^{b}|{\simeq}\frac{3^{a}}{\sqrt{4\pi a/3}}\exp\left(-\frac{3b^{2}}{4a}\right).$ (126)
Hence, in the asymptotic limit $|\mathcal{D}_{L}^{0}|$ takes the form:
$|\mathcal{D}_{L}^{0}|{=}\frac{3^{L{+}1/2}}{2\sqrt{\pi L}},$ (127)
@@
$|\mathbb{V}_{n}\rangle{=}\frac{1}{\sqrt{|\mathcal{D}_{L}^{M_{n}/2}|}}\sum_{k{=}K_{\rm min}}^{K_{\rm max}}$
@@
$|\mathbb{V}_{n}\rangle{=}$ $\sum_{k{=}K_{\rm min}}^{K_{\rm max}}\Bigg[\frac{\sqrt{|\mathcal{D}_{L/2}^{k}|}\sqrt{|\mathcal{D}_{L/2}^{M_{n}/2-k}|}}{\sqrt{|\mathcal{D}_{L}^{M_{n}/2}|}}$ (130)
$\left(\sum_{|\vec{b}\rangle{\in}\mathcal{D}_{L/2}^{k}}\frac{1}{\sqrt{|\mathcal{D}_{L/2}^{k}|}}|\vec{b},\vec{b}\rangle_{A_{f}}\right){\otimes}$
$\left(\frac{1}{\sqrt{|\mathcal{D}_{L/2}^{M_{n}/2-k}|}}\sum_{|\vec{b^{\prime}}\rangle{\in}\mathcal{D}_{L/2}^{M_{n}/2-k}}|\vec{b^{\prime}},\vec{b^{\prime}}\rangle_{B_{f}}\right)\Bigg]$ (131)
@@
$\lambda_{k}$ ${=}\frac{\sqrt{|\mathcal{D}_{L/2}^{k}|}\sqrt{|\mathcal{D}_{L/2}^{M_{n}/2-k}|}}{\sqrt{|\mathcal{D}_{L}^{M_{n}/2}|}}.$ (132)
@@
$\lambda_{k}^{2}{=}\frac{|\mathcal{D}_{L/2}^{k}|^{2}}{|\mathcal{D}_{L}^{0}|}~~~{\rm as}~~~|\mathcal{D}_{L/2}^{k}|{=}|\mathcal{D}_{L/2}^{-k}|.$ (133)
From Eq. (126), substituting the asymptotic forms of the trinomial coefficients $|\mathcal{D}_{a}^{b}|$ we find

### v2

@@ IV.4 Additional family of scars: Mirror-dimer states with unconstrained spins
where each term couples two nearest-neighbors that are mirror images with respect to the chosen reflection axis. Again the Hamiltonian $H_{XY}{\equiv}\sum_{j{=}1}^{L}\mathcal{M}_{j}^{k}$. We find that $H_{XY}$ possesses a set of simultaneous zero modes of all the $\mathcal{M}_{j}^{k}$, which we denote by $|\mathbb{M}_{m,m^{\prime}}^{k}\rangle$ (see Appendix C.4):
@@
where $m,m^{\prime}{\in}\{-1,0,1\}$ and $|\phi^{\pm}\rangle$ are maximally entangled spin-$1$ dimers defined previously in Eq. (32). Each state $|\mathbb{M}_{m,m^{\prime}}^{k}\rangle$ therefore comprises a fully dimerized background that is symmetric about the reflection axis passing through $k$ and k+Lk{+}L, along with two freely configurable spins at the ends of this axis in the state $|m,m^{\prime}\rangle$, which can take any of the nine possible two-site spin-$1$ configurations. A schematic of $|\mathbb{M}_{m,m^{\prime}}^{1}\rangle$ is shown in Fig. 12(b). The total magnetization of the state $|\mathbb{M}_{m,m^{\prime}}^{k}\rangle$ is $M_{m,m^{\prime}}{=}m{+}m^{\prime}$, allowing these states to appear in both even- and odd-magnetization sectors. This does not contradict the earlier chiral-symmetry arguments, which showed that the number of zero-energy eigenstates is lower bounded by zero in the odd-magnetization sectors [see Eq. (6)]; states can certainly exist even if this lower bound is zero. A representative state $|\mathbb{M}_{m,m^{\prime}}^{1}\rangle$ for 2L=122L{=}12 is shown in Fig. 12(b).
Furthermore, these mirror-dimer states $|\mathbb{M}_{m,m^{\prime}}^{k}\rangle$ are also annihilated by the NNNN $XY$ interaction term $V_{\rm NNNN}$ defined in Eq. (26), making them exact eigenstates of a fully non-integrable Hamiltonian $H_{\rm mirror}{=}H_{0}{+}V_{\rm NNNN}$ with energy determined by magnetization as $E_{m,m^{\prime}}^{k}{=}h(m{+}m^{\prime})$. The entanglement properties of these states depend strongly on the bipartition, as expected from their structure. If the subsystem lies entirely on one side of the reflection axis, i.e., contains exactly one site from each entangled dimer, the EE grows linearly with the subsystem size, displaying volume-law scaling as each dimer cut by the bipartition contributes a fixed entropy of $\ln(3)$. In contrast, if the subsystem includes only mirror-reflected site pairs, i.e., both sites of the dimers reside in the same subsystem, the state factorizes, and its entanglement entropy vanishes.
@@ V Conclusion and outlook
Partly motivated by recent works on identifying non-thermal eigenstates from quantum interference effects in Fock space, i.e., so-called Fock-Space Cage (FSC) states [69, 70, 71], we studied the Fock-space graph of this model to uncover an entire hierarchy of FSC eigenstates, parameterized by a positive integer separation $r$ (Sec. III), each realizing a distinct interference-protected cage structure within the zero-energy manifold. We used this perspective to also show that previously known scars [31, 32] also admit a natural interpretation as FSC states, where destructive interference confines their support to closed loops in the Fock-space graph. These novel quasiparticle towers of eigenstates we found are also isolated from the thermal continuum in entanglement and exhibit subextensive entanglement entropy, appearing as clear outliers well below the Page value. Coherent superpositions of these ladder states, while not simple product states but possessing finite, low entanglement even in the thermodynamic limit, display long-lived fidelity oscillations, the defining dynamical hallmark of scarring.
@@
The lowest members of these families are simple-looking on the Fock-space graph and admit transparent geometric interpretations based on simple pairwise interference cancellations, which qualify them as FSC states. However, for higher members, the interference patterns become increasingly complex and multipath in nature, making the geometric picture less intuitive and points toward a deeper organizing principle. To uncover this hidden principle, we turned to ideas from the commutant algebra framework (Sec. IV), which motivates the understanding of QMBS as simultaneous eigenstates of multiple non-commuting local operators.
@@
By systematically reorganizing the terms in the spin-$1$ XYXY Hamiltonian into carefully chosen families of non-commuting operators—constructed from periodically clustered, antipodal, and mirror-symmetric combinations of exchange bonds—we identified the algebraic structure that leads to the existence of the FSC states. In fact, all the FSC states we identified can be reinterpreted in the commutant algebra framework as simultaneous eigenstates of some set of local operators (i.e., singlets of some bond algebras), but such singlets of bond algebras do not always lend themselves to a simple interpretation in terms of FSC states. Using the commutant framework, we also uncovered two additional classes of exact eigenstates. The first of these, arising from antipodal magnon pairings, displays a striking bipartition-dependent entanglement structure: while exhibiting volume-law scaling under standard cuts, they show sub-extensive EE under fine-tuned antipodal partitions. The second family, consisting of mirror-dimer states, features reflection-symmetric dimer configurations with two unconstrained spins. These unconstrained spins are completely free—any choice of their local configuration yields an exact zero-energy eigenstate of $H_{XY}$, which hints at their potential applications in quantum information processing. Moreover, the mirror-dimer and volume-entangled states identified here consist of long-range entangled spin pairs, which are structurally similar to nonlocal Bell-pair states, for which a concrete protocol for realization in programmable quantum simulators has been recently proposed [96]. In addition, a scheme to realize volume-entangled exact scar states in the PXP model on near-term Rydberg quantum simulators was proposed in Ref. [21]. These works suggest that the mirror-dimer and volume-entangled states can potentially be prepared on such platforms, and it would be interesting to come up with concrete methods to realize them experimentally in future.
@@
A concise summary of all exact scar families, both previously known and newly discovered, is presented in Table 2. The table lists their total magnetizations, minimal generating operator sets within the commutant algebra framework, and characteristic entanglement properties, serving as a unified map of how many kinds of nonthermal manifolds of states — from frustration-free bimagnons to highly entangled mirror-dimer states—emerge from common algebraic principles. In the future, it would be important to uncover other associated properties of these states, such as their stability under weak generic perturbations [97, 98], the associated asymptotic quantum many-body scars [99, 100, 101, 102, 103], and potentially a complete classification of the distinct kinds of local parent Hamiltonians that could realize these as eigenstates [34, 87, 74, 104]. Finally, it would also be interesting to explore whether there are yet additional QMBS hiding in the spin-$1$ XYXY model that might be revealed as simultaneous eigenstates of other sets of non-commuting operators that we have not considered in this work.
@@ Acknowledgements.
We acknowledge useful discussions with Nicholas O’Dea, Frank Pollmann, Arnab Sen, and Maksym Serbyn. This work was undertaken on the Nandadevi supercomputer, which is maintained and supported by the High-Performance Computing Center at the Institute of Mathematical Sciences, India. S. Mohapatra is grateful to the School of Natural Sciences at the Technical University of Munich (TUM) for its hospitality in the summer of 2025. S. Moudgalya acknowledges support from the Munich Center for Quantum Science and Technology (MCQST) and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC–2111–390814868.
@@ C.3.1 Absence of a single raising operator for the volume-entangled tower
$|\psi_{\rm vol}^{\rm init}\rangle{=}\frac{1}{\sqrt{\mathcal{D}_{L}}}\sum_{|\vec{b^{\prime}}\rangle{\in}\mathcal{D}_{L}}\hat{\mathcal{C}}|\vec{b^{\prime}}\rangle_{1,{\cdots},L}{\otimes}|\vec{b^{\prime}}\rangle_{L{+}1,{\cdots},2L},$ (124)
@@
$|\mathbb{V}_{n}\rangle{=}\frac{1}{\sqrt{\mathcal{D}_{L}^{M_{n}/2}}}\sum_{|\vec{b}\rangle{\in}\mathcal{D}_{L}^{M_{n}/2}}\hat{\mathcal{C}}|\vec{b}\rangle_{1,{\cdots},L}{\otimes}|\vec{b}\rangle_{L{+}1,{\cdots},2L},$ (125)
where $\vec{b}$ runs over all spin-$1$ basis configurations with magnetization $M_{n}/2$ and $\mathcal{D}_{L}^{M_{n}/2}$ denote this Hilbert space dimension. Eq. (125) represents the Schmidt decomposition of $|\mathbb{V}_{n}\rangle$ for the standard bipartition $A_{s}{=}\{1,{\cdots},L\},~~B_{s}{=}\{L{+}1,{\cdots},2L\}$) with $\mathcal{D}_{L}^{M_{n}/2}$, with all non-zero Schmidt coefficients, $\lambda_{k}$, equal to $1/{\sqrt{\mathcal{D}_{L}^{M_{n}/2}}}$. Thus, the von Neumann EE of $|\mathbb{V}_{n}\rangle$ for this bipartition is
$\mathcal{S}_{A_{s}}(n){=}\sum_{k}\lambda_{k}^{2}\ln\lambda_{k}^{2}{=}\ln(\mathcal{D}_{L}^{M_{n}/2}).$ (126)
We now evaluate the $\mathcal{S}_{A_{s}}(n{=}L)$ for the mid-spectrum state $|\mathbb{V}_{n{=}L}\rangle$ (lying in the magnetization sector $M_{L}{=}0)$ in the large $L$ limit. We recall that the trinomial coefficient $\mathcal{D}_{a}^{b}$ (as defined in Eq. (5)) represents the coefficient of $x^{b}$ in the expansion of $(1{+}x{+}x^{{-}1})^{a}$. For large values of $a$, following a similar derivation as done in App. A, we get the approximate Gaussian form of $\mathcal{D}_{a}^{b}$:
$\mathcal{D}_{a}^{b}{\simeq}\frac{3^{a}}{\sqrt{4\pi a/3}}\exp\left(-\frac{3b^{2}}{4a}\right).$ (127)
Hence, in the asymptotic limit $\mathcal{D}_{L}^{0}$ takes the form:
$\mathcal{D}_{L}^{0}{=}\frac{3^{L{+}1/2}}{2\sqrt{\pi L}},$ (128)
@@
$|\mathbb{V}_{n}\rangle{=}\frac{1}{\sqrt{\mathcal{D}_{L}^{M_{n}/2}}}\sum_{k{=}K_{\rm min}}^{K_{\rm max}}$
@@
$|\mathbb{V}_{n}\rangle{=}$ $\sum_{k{=}K_{\rm min}}^{K_{\rm max}}\Bigg[\frac{\sqrt{\mathcal{D}_{L/2}^{k}}\sqrt{\mathcal{D}_{L/2}^{M_{n}/2-k}}}{\sqrt{\mathcal{D}_{L}^{M_{n}/2}}}$ (131)
$\left(\sum_{|\vec{b}\rangle{\in}\mathcal{D}_{L/2}^{k}}\frac{1}{\sqrt{\mathcal{D}_{L/2}^{k}}}|\vec{b},\vec{b}\rangle_{A_{f}}\right){\otimes}$
$\left(\frac{1}{\sqrt{\mathcal{D}_{L/2}^{M_{n}/2-k}}}\sum_{|\vec{b^{\prime}}\rangle{\in}\mathcal{D}_{L/2}^{M_{n}/2-k}}|\vec{b^{\prime}},\vec{b^{\prime}}\rangle_{B_{f}}\right)\Bigg]$ (132)
@@
$\lambda_{k}$ ${=}\frac{\sqrt{\mathcal{D}_{L/2}^{k}}\sqrt{\mathcal{D}_{L/2}^{M_{n}/2-k}}}{\sqrt{\mathcal{D}_{L}^{M_{n}/2}}}.$ (133)
@@
$\lambda_{k}^{2}{=}\frac{(\mathcal{D}_{L/2}^{k})^{2}}{\mathcal{D}_{L}^{0}}~~~{\rm as}~~~\mathcal{D}_{L/2}^{k}{=}\mathcal{D}_{L/2}^{-k}.$ (134)
From Eq. (127), substituting the asymptotic forms of the trinomial coefficients $\mathcal{D}_{a}^{b}$ we find